OpenStudy (anonymous):

what is -(14x)^0y^-7z in simplified form?

OpenStudy (anonymous):

start with the fact that \(b^0=1\) and get \[-1\times y^{-7}\times z\]

OpenStudy (anonymous):

then use \(b^{-n}=\frac{1}{b^n}\) to get \[-\frac{z}{y^7}\]

OpenStudy (anonymous):

Thanx so much